OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths place if necessary. 11x2 - x - 3 = 0 A. {-1, 1.2} B. {-11, 13.2} C. {-0.5, 0.6} D. no real solution

OpenStudy (anonymous):

you have a equation in the form ax^2 + bx + c so a=11 b=-1 c=-3 plug these values into the formula \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[ b^2 -4 a c =133>0 \] We have two real roots. Find them using the formula that @Tyler1992 gave you above.

OpenStudy (anonymous):

\[-(-1)\pm \sqrt{-1 -4(11)(-3)}\]

OpenStudy (anonymous):

over 2(11)

OpenStudy (anonymous):

good now you just have to simplify and you're done

OpenStudy (anonymous):

\[\frac{-(-1) \pm \sqrt{(-1)^2 - 4(11)(-3)}}{2(11)} = \frac{1 \pm \sqrt{133}}{22}\]

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

0.5,0.6

OpenStudy (anonymous):

the answers should be -5, and 6 not .5 and .6 so im guessing its a typo in the choices